An experiment involves 17 participants. From these, a group of 5 participants is to be tested under a special condition. How many groups of 5 participants can be chosen, assuming that the order in which the participants are chosen is irrelevant?

Answer by Mathtut(3670)

17*16*15*14*13/5*4*3*2=742560/120=6188